Solve by factoring. Show work
2k^2+11k=-5

Answer by Alan3354(69443)

Solve by factoring. Show work
2k^2+11k=-5
-------
2k^2 + 11k + 5 = 0
(2k + 1)*(k + 5) = 0
k = -5
k = -1/2